OpenStudy (anonymous):

you have 20 shirts in your closet. 4 blue, 7 red, 3 green and 6 black. you pick one without looking. what is the probability that you pick a blue or black shirt(1/4, 1/2, 3/50)?

OpenStudy (dumbcow):

there are 20 total shirts there are 10 blue or black shirts --> P = 10/20 = 1/2